OpenStudy (anonymous):

If f(x) = 3x + 2 and g(x) = 2x – 2, what is (f – g)(x)? x + 4 x – 2 x 5x – 2 x – 4

OpenStudy (anonymous):

put a minus sign between them use parentheses

OpenStudy (anonymous):

what are you talking about

OpenStudy (anonymous):

\[-\] is a minus sign

OpenStudy (anonymous):

\[f-g=(3x+2)-(2x-2)\]

OpenStudy (anonymous):

aka put a minus sign between them use parentheses

OpenStudy (♪chibiterasu):

Ok, the site went down while I was answering this. "We can combine the functions based on the question. (f-g)(x) 3x + 2 - 2x - 2 We can combine like terms. Do you know how to do that?"

OpenStudy (anonymous):

i know what it is ill get some else help if your gonna be smart about it

OpenStudy (anonymous):

no i do not know how to do that

OpenStudy (♪chibiterasu):

Ok, well. If you had 4x and -5x you could combine them. They are like terms because they both have the variable 'x'. Since -5x is negative, we have to subtract 4x - 5x = -1x or just -x. Now if we had 2x and 6x^2, we can't combine them. They both have the variable 'x' but 6x^2 has an exponent, so they are not the same and we cannot combine.

OpenStudy (♪chibiterasu):

Our like terms in this problem are: 3x and -2x 2 and -2

OpenStudy (anonymous):

so my answer is x

OpenStudy (♪chibiterasu):

Yeah.

OpenStudy (♪chibiterasu):

3x - 2x = x 2 - 2 = 0 x :3

OpenStudy (anonymous):

okay i got three more questions

OpenStudy (♪chibiterasu):

I have to go now, sorry! D:

OpenStudy (anonymous):

too bad that answer is incorrect

OpenStudy (anonymous):

its okay thanks for the help

OpenStudy (anonymous):

Andy is wrapping gifts for his family. He ties 25 inches of ribbon around each gift. Which function gives the relationship between the number of gifts (g) and the length of ribbon used (r)? The function representing the relationship between the length of ribbon and the number of gifts is

OpenStudy (anonymous):

lets do the first one correctly

OpenStudy (anonymous):

ive already answered it

OpenStudy (anonymous):

what did you put?

OpenStudy (anonymous):

i can use help on the last few tho

OpenStudy (anonymous):

and i put x

OpenStudy (anonymous):

it is wrong

OpenStudy (anonymous):

i have a few more to go if i get right i should pass

OpenStudy (anonymous):

going to next class give me like 10 minutes can you help with this

OpenStudy (anonymous):

Andy is wrapping gifts for his family. He ties 25 inches of ribbon around each gift. Which function gives the relationship between the number of gifts (g) and the length of ribbon used (r)? The function representing the relationship between the length of ribbon and the number of gifts is